OpenStudy (anonymous):

What is the speed of an electron that has been accelerated from rest through a potential difference of 1020 V?

OpenStudy (jamesj):

E = F/q hence the work divided by q, \[ \frac{W}{q} = \frac{Fd}{q} = Ed \] and \( Ed = V \), the potential difference. Therefore the work done \[ W = Vq \] So you can calculate the work done on the electron. By conservation of energy, that work is now somewhere, and that somewhere is all kinetic energy. Provided the velocity isn't relativistic, you can calculate it as being equal to \( KE = \frac{1}{2}mv^2 \)

OpenStudy (anonymous):

In this problem I'm solving for the velocity because the the final units are m/s. So, how would you find KE to that I could solve for v in the last equation?

OpenStudy (aravindg):

its KE is 850 eV

OpenStudy (aravindg):

convert this into joulses and equate to 1/2mv^2

OpenStudy (aravindg):

and get ur v!

OpenStudy (anonymous):

I hat to say it but i still don't get it. Do I do anything with the 1020V? What i did you originally wa Vq=1/2mv^2 but it said i didnt get the right answer. Am i suppose to do something like that? If not, please try to explain what you all are saying?

OpenStudy (jamesj):

The work done on a charge q moving through a potential of V is W = Vq. Here, the charge on an electron is e = 1.602 x 10^-19 C and therefore the work done on it is W = Vq = (1020)(1.602 x 10^-19) Joules Now equate that to the change in KE, which, as as the electron started at rest is just the ending kinetic energy, \( \frac{1}{2}mv^2 \). Make sense?

OpenStudy (anonymous):

Yes thank you so much